# A Stochastic–Geometrical Framework for Object Pose Estimation based on Mixture Models Avoiding the Correspondence Problem

### v1

@@ Abstract
Background: Pose estimation of rigid objects is a practical challenge in optical metrology and computer vision. In this paper a novel stochastic–geometrical modeling framework for object pose estimation is presented based on observing multiple feature points.
Methods: This stochastic–geometrical framework utilizes mixture models for the feature point densities in object space as well as for interpreting real measurements. Direct advantages of this approach are the avoidance to resolve individual feature correspondences and to incorporate correct stochastic dependencies in multi–view applications. First, the general modeling framework is presented, second, a general algorithm for pose estimation is derived, and third, two example models for a camera setup as well as a lateration setup are presented.
Results: The numerical experiments show the effectiveness of this modeling and general algorithm by investigating four simulation scenarios for three different observation systems, including the dependence on measurement resolution, object deformations as well as strong measurement noise. It can be concluded that the probabilistic modeling of pose estimation based on mixture models can lead to accurate and robust pose estimations.
@@ 1 Introduction
Pose estimation is a widely applied task in computer vision since the 1980’s by Fischler et al.[10] and other disciplines that require, for example, the automatic tracking of objects or single– or multiple–view projection based estimation, i.e. see Groschlüteret al.[11], He et al.[13] and Sahin et al.[20] for recent detailed reviews of state–of–the–art techniques. The goal is always to find the position and orientation of a predefined 3D object, which is observed typically by cameras. To achieve this algorithmically, many advanced ad–hoc loss functions and dedicated algorithms are applied successfully, most recently with the use of convolutional neural networks (CNNs) (e.g. Kendall et al.[16], Bui et al.[2], Li et al.[17], and Wang et al.[24]) or kernel–based algorithms (e.g. Haopeng et al.[12] and Nie et al.[18]). The exemplary background of this work is the rigid body pose estimation in the industrial application of tracking of laser scanning devices in optical metrology for which typically model based approaches, as presented in this paper, are state–of–the–art. The challenge for such devices is to find a balance between high accuracy and robust rigid body pose estimation, maintainability/servicability, safety and time performance regarding application and calibration of these products. Useful approaches to a more general perspective on object pose estimation are modeling frameworks, which have been presented with the purpose to subdivide the process of pose estimation for multiple objects in a scene into smaller, modular steps and provide a review of possible specific algorithms of classical image processing, feature segmentation, image and 3D model matching and pose optimization (cp. Collet et al.[5]) or different architectures of CNNs for this purpose (cp. Li et al.[17]). As demonstrated recently by He et al.[13] major challenges of 6D pose estimation to this day are the robustness of algorithms relative to image distortions and the problem of how to satisfyingly deal with deformed objects.
@@
The closest applications of such a stochastic/probabilistic viewpoint to pose estimation has been applied by Erkent et al.[9], Teney et al.[23] and Teney et al.[22] with similarities especially by interpreting the observational measurement as a mixture model utilizing Gaussian kernels. Unfortunately only a partial stochastic model was applied which in turn led to ad–hoc mathematical arguments and results, e.g. utilizing products of cross correlations in multi–view applications without strict derivations. Further, also the concept of modeling predicted marginal object point Gaussian Mixture Models directly in measurement space for 6D pose estimation has been introduced in Brachmann et al.[3] in order to increase the robustness of estimation. Although these approaches show high performance in their presented applications and prove the utility of stochastic models, in this paper the point of view is taken that the full potential of stochastic modeling for pose estimation in a general observation scenario is not addressed, since e.g. it is avoided to apply a clear and straightforward argumentation by random variables (including their dependency structure especially for multi–view applications), the applied density function types are limited to specific classes and the presented setting is limited to camera observations only. It is also noted that the stochastic modeling of feature points in object space (again limited to specific density functions, such as Gaussian Mixture Models) is currently successfully applied in point set registration, e.g. in Chen et al.[4] and Deng et al.[6]. Preliminary work for a specific example for a strictly stochastic framework for tumor pose estimation was presented in the field of medical imaging for radiotherapy with the purpose of optimal patient positioning by the author (Hoegele et al.[14, 15]).
@@ 2.1.2 Modeling the Observation Prediction from Object Space to Measurement Space
In general, it is not necessary to work with individual feature points $\vec{z}_{i}$ ($i=1,..,N$) to construct such a mixture model, but for the simplicity of the argumentation this is used throughout the paper. For example, also line segments or other structures representing observable parts of the object and their uncertainties can be utilized in the construction of this mixture model.
@@ 2.2.1 Modeling the Object Pose Estimation with a Single Observation
It is a powerful extension of this framework interpreting the density functions of feature coordinates of the object $\vec{Z}_{i}\sim q_{i}(\vec{z})$ not only as the unknown position of a truly rigid body but also of a deformed object. Of course, this stochastic formulation of object pose estimation is then still straightforwardly applied without change for this much more difficult task since it is only a change in the interpretation of these density functions. For example, if the $6$D pose of a rigid body needs to be determined, it is practically possible to incorporate the probable uncertainties of the object form due to gravitational influences into the pose estimation by updating these density functions.
@@ 2.2.2 Extension to Multiple Observations with known Relative Geometry
This object pose estimation can be easily extended to multiple observations if the relative geometry of the observations are fixed and known in advance, e.g. by a static multiple camera device. The purpose again, is to find the parameters $\vec{\beta}$ describing the status of the object relative to the first observation geometry as in the previous derivations.
@@
A common misconception is that the matching functions between observation and prediction for different observations of the same object are regarded as independent. In consequence, these matching functions of different observations $l=1,..,L$ are simply multiplied. It is pointed out that this is wrong, which can be realized by working directly with the difference random variables $\vec{Y}_{l}-\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)$ ($l=1,..,L$). These difference random variables are not independent for different observations $l$, since the prediction part $\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)$ depends on the commonly observed object features described by $\vec{Z}$. On the other side, the individual measurement random variables $\vec{Y}_{l}$ can safely be assumed to be independent, especially for electronic measurement devices which have no memory in their sensors. To illustrate the meaning of this argument, it is emphasized that the obvious multiplication of matching functions of each individual observation misses the correct dependency structure, e.g.
$f_{\bigcap\limits_{l=1}^{L}\vec{Y}_{l}-\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)}(\vec{0})\,\neq\,\prod\limits_{l=1}^{L}\,f_{\vec{Y}_{l}-\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)}(\vec{0})\;.$
For later comparison of the formula, following this wrong argumentation one would arrive straightforwardly at the implementation formula via Monte–Carlo approximation
On the contrary, we have to introduce the following argumentation if we consider the correct dependency structure of the random variables
@@
again with $\vec{z}_{i,r}$ ($r=1,..,R$) being random samples of the pdf $q_{i}(\vec{z})$ ($i=1,..,N$). It is pointed to the difference of this optimization compared to the optimization when neglecting the dependency structure of the random variables, which constitutes itself by the different position of the product. It is probable that utilizing a wrong assumption, such as the independency of matching functions in a multi–view application, leads to inferior results.
@@ 2.3.2 Camera System with Feature Coordinate Densities and Parallel Beam Projection Geometry
In this application, we are observing a $3$D object with feature coordinates $\vec{z}_{i}\in\mathbb{R}^{3}$, ($i=1,..,N$), with $2$D camera observations/projections. We assume a low knowledge level regarding the feature coordinates, specifically, that they are symmetrically normal distributed, i.e. $q_{i}(\vec{z})=\mathcal{N}(\vec{z}_{i},\text{diag}(\sigma_{i,1}^{2},\sigma_{i,2}^{2},\sigma_{i,3}^{2}))$.
Further, we assume a specific projection geometry: Parallel beam projection from 3D-object to 2D-camera while the 3D-object can be shifted and rotated only around the $z_{3}$-axis:
$\vec{p}(\vec{z};(\varphi,\vec{w}),\gamma)$
$:=\left(\begin{array}[]{ccc}\cos(\varphi+\gamma)&\sin(\varphi+\gamma)&0\\ 0&0&1\end{array}\right)\cdot\left(\vec{z}+\vec{w}\right)=\left(\begin{array}[]{c}\cos(\varphi+\gamma)\cdot(z_{1}+w_{1})+\sin(\varphi+\gamma)\cdot(z_{2}+w_{2})\\ z_{3}+w_{3}\end{array}\right)\;.$
The parameter $\gamma$ describes in this case how the observation geometry is rotated around the object (or more general what the zero angle of rotation in this specific observation is).
@@
$\text{arg max}_{\,(\varphi,\vec{w})}\;\frac{1}{N\,R}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}g\left(\begin{array}[]{c}\cos(\varphi+\gamma)\cdot(z_{i,r,1}+w_{1})+\sin(\varphi+\gamma)\cdot(z_{i,r,2}+w_{2})\\ z_{i,r,3}+w_{3}\end{array}\right)\;,$
@@
For the extension to multiple projections, we present the formula for pose estimation with two orthogonal projections, which means that we have $\gamma_{1}$ as initial angle of the first observation and $\gamma_{2}=\gamma_{1}+\tfrac{\pi}{2}$ in the second observation. This can be summarized by the specific application of formula (4):
$\text{arg max}_{\,(\varphi,\vec{w})}\;\frac{1}{N}\sum\limits_{i=1}^{N}\iiint_{\mathbb{R}^{3}}g_{1}\left(\vec{p}\left(\vec{z};(\varphi,\vec{w}),\gamma_{1}\right)\right)\cdot g_{2}\left(\vec{p}\left(\vec{z};(\varphi,\vec{w}),\gamma_{1}+\tfrac{\pi}{2}\right)\right)\cdot q_{i}(\vec{z})\;\text{d}\vec{z}\;,$
@@
$\text{arg max}_{\,(\varphi,\vec{w})}\;\frac{1}{NR}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}g_{1}\left(\vec{p}\left(\vec{z}_{i,r};(\varphi,\vec{w}),\gamma_{1}\right)\right)\cdot g_{2}\left(\vec{p}\left(\vec{z}_{i,r};(\varphi,\vec{w}),\gamma_{1}+\tfrac{\pi}{2}\right)\right)\;,$
@@
A short comment regarding the calibration of this camera system: It has to be noted that for the calibration of this specific geometry $\gamma_{1}$ can be set arbitrarily, since the only relevant information is that $\gamma_{2}:=\gamma_{1}+\tfrac{\pi}{2}$. In practice one would (if there are no arguments against that) set $\gamma_{1}:=0$ for the calibration and let $\gamma_{2}$ optimize in order to find the deviations from the relative observation angle $\tfrac{\pi}{2}$.
@@ 2.3.3 Lateration System with Feature Coordinate Densities and Rigid Affine Transformation Geometry
As second example, we are focusing on a lateration system, which measures the distances from a lateration point to each feature, such as radar or lidar devices. This naturally leads to $1$D observations of the $3$D objection consisting of the feature coordinates $\vec{z}_{i}$ ($i=1,..,N$). This time we assume to not know the position of the feature coordinates exactly, and, in consequence, utilize further unspecified density functions $q_{i}(\vec{z})$.
The projection function in this application consists of two steps. First, the coordinate transformation of the object, and second, the projection to distances. We get with $\vec{\beta}:=(V,\vec{w})$
@@
$p(\vec{z};(V,\vec{w}),\vec{\gamma})$ $:=\sqrt{\left(u_{1}\left(\vec{z};(V,\vec{w})\right)-\gamma_{1}\right)^{2}+\left(u_{2}\left(\vec{z};(V,\vec{w})\right)-\gamma_{2}\right)^{2}+\left(u_{3}\left(\vec{z};(V,\vec{w})\right)-\gamma_{3}\right)^{2}}\;.$
With this example, it is demonstrated that the argument presented in this paper does not depend on well–behaved pdfs. The objective function for one projection stays general and we arrive at the formula including the Monte–Carlo approximation as specific application of formula (2):
$\text{arg max}_{\,(V,\vec{w})}\;\frac{1}{N\,R}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}g\left({p}\left(\vec{z}_{i,r};(V,\vec{w}),\vec{\gamma}\right)\right)\;,\text{ subject to }V\text{ with orthonormal columns},$
with $p$ being the described rigid affine transformation and projection, and $\vec{\gamma}$ the position of the lateration source in $3$D. It has to be mentioned that in this general formulation the specific orthogonality of $V$ directly leads to constraints in optimization, such as a system of nonlinear equations for the entries of $V$, which would need to be dealt with in the optimization routine. Of course, orthogonality of $V$ can be also enforced by the construction of the matrix, such as a rotation matrix and utilize the angles as degrees of freedom as demonstrated in the results section 3.3.
@@
$\text{arg max}_{\,(V,\vec{w})}\;\frac{1}{N\,R}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}\prod\limits_{l=1}^{3}\,g_{l}\left({p}\left(\vec{z}_{i,r};(V,\vec{w}),\vec{\gamma}_{l}\right)\right)\;,\text{ subject to }V\text{ with orthonormal columns},$
with $\vec{z}_{i,r}$ sampled from the densities $q_{i}(\vec{z})$.
A short comment regarding the calibration of the lateration system: Three lateration devices have specific coordinates $\vec{\gamma}_{1},\vec{\gamma}_{2},\vec{\gamma}_{3}$. Since the absolute position of the system in a global coordinate system is not important for practical measurements with that system, one can set $\vec{\gamma}_{1}=(0,0,0)^{T}$, $\vec{\gamma}_{2}=(c_{2,1},0,0)^{T}$ and $\vec{\gamma}_{3}=(c_{3,1},c_{3,2},0)^{T}$ and, consequently, obtain pose estimations relative to this defined geometry.

### v4

@@ Abstract
Background: Pose estimation of rigid objects is a practical challenge in optical metrology and computer vision. This paper presents a novel stochastic–geometrical modeling framework for object pose estimation based on observing multiple feature points.
Methods: This framework utilizes mixture models for feature point densities in object space and for interpreting real measurements. Advantages are the avoidance to resolve individual feature correspondences and to incorporate correct stochastic dependencies in multi–view applications. First, the general modeling framework is presented, second, a general algorithm for pose estimation is derived, and third, two example models (camera and lateration setup) are presented.
Results: Numerical experiments show the effectiveness of this modeling and general algorithm by presenting four simulation scenarios for three observation systems, including the dependence on measurement resolution, object deformations and measurement noise. Probabilistic modeling utilizing mixture models shows the potential for accurate and robust pose estimations while avoiding the correspondence problem.
@@ 1 Introduction
Pose estimation is a widely applied task in computer vision since the 1980’s by Fischler et al.[10] and other disciplines that require, for example, the automatic tracking of objects or single– or multiple–view projection based estimation, i.e. see Groschlüteret al.[11], He et al.[13] and Sahin et al.[20] for recent detailed reviews of state–of–the–art techniques. The goal is always to find the position and orientation of a predefined 3D object, which is observed typically by cameras. To achieve this algorithmically, many advanced ad–hoc loss functions and dedicated algorithms are applied successfully, most recently with the use of convolutional neural networks (CNNs) (e.g. Kendall et al.[16], Bui et al.[2], Li et al.[17], and Wang et al.[24]) or kernel–based algorithms (e.g. Haopeng et al.[12] and Nie et al.[18]). The exemplary background of this work is the rigid body pose estimation in the industrial application of tracking of laser scanning devices in optical metrology for which typically model based approaches, as presented in this paper, are state–of–the–art. The challenge for such devices is to find a balance between high accuracy and robust rigid body pose estimation, maintainability/servicability, safety and time performance regarding application and calibration of these products. Useful approaches to a more general perspective on object pose estimation are modeling frameworks, which have been presented with the purpose to subdivide the process of pose estimation for multiple objects in a scene into smaller, modular steps and provide a review of possible specific algorithms of classical image processing, feature segmentation, image and 3D model matching and pose optimization (cp. Collet et al.[5]) or different architectures of CNNs for this purpose (cp. Li et al.[17]). As demonstrated recently by He et al.[13], the major challenges of 6D pose estimation to this day are the robustness of algorithms relative to image distortions and the problem of how to satisfyingly deal with deformed objects.
@@
The closest applications of such a stochastic/probabilistic viewpoint to pose estimation has been applied by Erkent et al.[9], Teney et al.[23] and Teney et al.[22] with similarities especially by interpreting the observational measurement as a mixture model utilizing Gaussian kernels. Unfortunately only a partial stochastic model was applied which in turn led to ad–hoc mathematical arguments and results, e.g. utilizing products of cross correlations in multi–view applications without strict derivations. Further, also the concept of modeling predicted marginal object point Gaussian Mixture Models directly in measurement space for 6D pose estimation has been introduced in Brachmann et al.[3] in order to increase the robustness of estimation. Although these approaches show high performance in their presented applications and prove the utility of stochastic models, in this paper the point of view is taken that the full potential of stochastic modeling for pose estimation in a general observation scenario is not addressed. For example, it is avoided to apply a clear and straightforward argumentation by random variables (including their dependency structure especially for multi–view applications), the applied density function types are limited to specific classes and the presented setting is limited to camera observations only. It is also noted that the stochastic modeling of feature points in object space (again limited to specific density functions, such as Gaussian Mixture Models) is currently successfully applied in point set registration, e.g. in Chen et al.[4] and Deng et al.[6]. Preliminary work for a specific example for a strictly stochastic framework for tumor pose estimation was presented in the field of medical imaging for radiotherapy with the purpose of optimal patient positioning by the author (Hoegele et al.[14, 15]).
@@ 2.1.2 Modeling the Observation Prediction from Object Space to Measurement Space
In general, it is not necessary to work with individual feature points $\vec{z}_{i}$ ($i=1,..,N$) to construct such a mixture model, but for the simplicity of the argumentation this is used throughout the paper. For example, also line segments (in order to approximate object edges) or other structures representing observable parts of the object and their uncertainties can be utilized in the construction of this mixture model. For example, other types of features could be surface structures of objects or even textures on an object surface. The crucial part will be that these features can be extracted well in the observations. For the practical implementation, the density function $q(\vec{z})$ needs to be modeled with high density areas at the location of these features.
@@ 2.2.1 Modeling the Object Pose Estimation with a Single Observation
It is a powerful extension of this framework interpreting the density functions of feature coordinates of the object $\vec{Z}_{i}\sim q_{i}(\vec{z})$ not only as the unknown position of a truly rigid body but also of a deformed object. Of course, this stochastic formulation of object pose estimation is then still straightforwardly applied without change for this much more difficult task since it is only a change in the interpretation of these density functions. For example, when tracking large and/or heavy rigid objects it is a practical concern that they slightly deform depending on the $6$D pose since the gravitational force is acting on the whole body. This may lead to individual feature displacements depending on the $6$D pose, which (moderately) violates the rigid body assumption. In the presented approach it is practically possible to incorporate such feature displacements as positional uncertainties into the pose estimation by updating the feature density functions with the range of individual feature displacements in the direction of the gravitational force.
@@ 2.2.2 Extension to Multiple Observations with known Relative Geometry
This object pose estimation can be easily extended to multiple observations if the relative geometry of the observations are fixed and known in advance, e.g. by a static multiple camera device. The purpose again, is to find the parameters $\vec{\beta}$ describing the status of the object relative to the observation geometry as in the previous derivations.
@@
Before we arrive at the main argument, we must clarify an important point: A common misconception is that the matching functions between observation and prediction for different observations of the same object are regarded as independent. In consequence, these matching functions of different observations $l=1,..,L$ are typically multiplied. It is pointed out that this is wrong, which can be realized by working directly with the difference random variables $\vec{Y}_{l}-\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)$ ($l=1,..,L$). These difference random variables are not independent for different observations $l$, since the prediction part $\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)$ depends on the commonly observed object features described by $\vec{Z}$. To illustrate the meaning of this argument, it is emphasized that the obvious multiplication of matching functions of each individual observation misses the correct dependency structure, e.g.
$f_{\bigcap\limits_{l=1}^{L}\vec{Y}_{l}-\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)}(\vec{0})\,\neq\,\prod\limits_{l=1}^{L}\,f_{\vec{Y}_{l}-\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)}(\vec{0})\;,$
with $f_{\bigcap\limits_{l=1}^{L}\vec{Y}_{l}-\vec{p}\left(\vec{Z};\vec{\beta},\vec{\gamma}_{l}\right)}$ being the joint probability distribution of the matching for all observations.
On the other side, the individual measurement random variables $\vec{Y}_{l}$ can safely be assumed to be independent, especially for electronic measurement devices which have no memory in their sensors.
To avoid this misconception, we have to introduce the following argumentation if we consider the correct dependency structure of the random variables
@@
again with $\vec{z}_{i,r}$ ($r=1,..,R$) being random samples of the pdf $q_{i}(\vec{z})$ ($i=1,..,N$).
For comparison of the formula, the result of the neglection of the dependency structure by multiplication of the matching function of different observations leads to the Monte–Carlo approximation
The difference constitutes itself by the different position of the product. It is probable that utilizing a wrong assumption, such as the independency of matching functions in a multi–view application, leads to inferior results.
@@ 2.3.2 Camera System with Feature Coordinate Densities and Parallel Beam Projection Geometry
In this example application, we are observing a $3$D object with feature coordinates $\vec{z}_{i}\in\mathbb{R}^{3}$, ($i=1,..,N$), with $2$D camera observations/projections. We assume a low knowledge level regarding the feature coordinates, specifically, that they are symmetrically normal distributed, i.e. $q_{i}(\vec{z})=\mathcal{N}(\vec{z}_{i},\text{diag}(\sigma_{i,1}^{2},\sigma_{i,2}^{2},\sigma_{i,3}^{2}))$.
Further, we assume a specific projection geometry: Parallel beam projection from 3D-object to 2D-camera while the 3D-object can be shifted diagonally and rotated only around the $z_{3}$-axis (this limited transformation will be used in the results subsection 3.2 for visualization purposes):
$\vec{p}_{1}(\vec{z};(\varphi,w))$ $:=\left(\begin{array}[]{c}\cos(\varphi)\cdot(z_{1}+w)+\sin(\varphi)\cdot(z_{2}+w)\\ -\sin(\varphi)\cdot(z_{1}+w)+\cos(\varphi)\cdot(z_{2}+w)\\ z_{3}+w\end{array}\right)\;,$
$\vec{p}_{2}(\vec{p}_{1}(\vec{z};(\varphi,w));\gamma)$ $:=\left(\begin{array}[]{ccc}\cos(\gamma)&\sin(\gamma)&0\\ 0&0&1\end{array}\right)\cdot\vec{p}_{1}(\vec{z};(\varphi,w))\;.$
The parameter $\gamma$ describes in this case how the cameras are rotated around the object. The overall measurement situation is illustrated in figure 1.
Figure 1: Schematic illustration of a camera system observation geometry and their connections to the mixture model approach.
@@
$\text{arg max}_{\,(\varphi,w)}\;\frac{1}{N\,R}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}g\left(\left(\begin{array}[]{ccc}\cos(\gamma)&\sin(\gamma)&0\\ 0&0&1\end{array}\right)\cdot\vec{p}_{1}(\vec{z}_{i,r};(\varphi,w))\right)\;,$
@@
For the extension to multiple projections, we present the formula for pose estimation with two orthogonal projections, which means that we have $\gamma_{1}$ as initial angle of the first observation and $\gamma_{2}:=\gamma_{1}+\tfrac{\pi}{2}$ in the second observation. This can be summarized by the specific application of formula (4):
$\text{arg max}_{\,(\varphi,w)}\;\frac{1}{N}\sum\limits_{i=1}^{N}\iiint\limits_{\mathbb{R}^{3}}g_{1}\left(\left(\begin{array}[]{ccc}\cos(\gamma_{1})&\sin(\gamma_{1})&0\\ 0&0&1\end{array}\right)\cdot\vec{p}_{1}(\vec{z};(\varphi,w))\right)\cdot\dots$
$g_{2}\left(\left(\begin{array}[]{ccc}\cos(\gamma_{1}+\frac{\pi}{2})&\sin(\gamma_{1}+\frac{\pi}{2})&0\\ 0&0&1\end{array}\right)\cdot\vec{p}_{1}(\vec{z};(\varphi,w))\right)\cdot q_{i}(\vec{z})\;\text{d}\vec{z}\;,$
@@
$\text{arg max}_{\,(\varphi,w)}\;\frac{1}{N\,R}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}\;g_{1}\left(\left(\begin{array}[]{ccc}\cos(\gamma_{1})&\sin(\gamma_{1})&0\\ 0&0&1\end{array}\right)\cdot\vec{p}_{1}(\vec{z}_{i,r};(\varphi,w))\right)\cdot\dots$
$g_{2}\left(\left(\begin{array}[]{ccc}\cos(\gamma_{1}+\frac{\pi}{2})&\sin(\gamma_{1}+\frac{\pi}{2})&0\\ 0&0&1\end{array}\right)\cdot\vec{p}_{1}(\vec{z}_{i,r};(\varphi,w))\right)\;,$
@@
A short comment regarding the calibration of this camera system: It has to be noted that for the calibration of this specific geometry parameter $\gamma_{1}$ can be set arbitrarily, since the only relevant information is that $\gamma_{2}:=\gamma_{1}+\tfrac{\pi}{2}$. In practice one would (if there are no arguments against that) set $\gamma_{1}:=0$ for the calibration and let $\gamma_{2}$ optimize in order to find the deviations from the relative observation angle $\tfrac{\pi}{2}$.
@@ 2.3.3 Lateration System with Feature Coordinate Densities and Rigid Affine Transformation Geometry
As second example, we are focusing on a ranging/lateration system, which measures the distances from a lateration point to each feature, such as radar or lidar devices. This naturally leads to $1$D observations of the $3$D objection consisting of the feature coordinates $\vec{z}_{i}$ ($i=1,..,N$). We assume to not know the position of the feature coordinates exactly, and, in consequence, utilize further unspecified density functions $q_{i}(\vec{z})$.
The projection function in this application consists of two steps. First, the coordinate transformation of the object as the general rigid affine transformation, and second, the projection to distances. We get with $\vec{\beta}:=(V,\vec{w})$ and $\vec{\gamma}$ the position of the lateration source in $3$D:
@@
$p_{2}(\vec{u};\vec{\gamma})$ $:=\sqrt{\left(u_{1}-\gamma_{1}\right)^{2}+\left(u_{2}-\gamma_{2}\right)^{2}+\left(u_{3}-\gamma_{3}\right)^{2}}\;.$
The overall measurement situation is demonstrated in figure 2.
Figure 2: Schematic illustration of a ranging system observation geometry and their connections to the mixture model approach.
With this example, it is demonstrated that the argument presented in this paper does neither depend on well–behaved pdfs in the modeling process (such as Gaussians) nor on a specific type of observation system. The objective function for one projection stays general and we arrive at the formula including the Monte–Carlo approximation as specific application of formula (2):
$\text{arg max}_{\,(V,\vec{w})}\;\frac{1}{N\,R}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}g\left(p_{2}\left(V^{T}\cdot(\vec{z}_{i,r}+\vec{w});\vec{\gamma}\right)\right)\;,$
subject to $V\in\mathbb{R}^{3\times 3}$ with orthonormal columns and with $\vec{z}_{i,r}$ sampled from the densities $q_{i}(\vec{z})$. It has to be mentioned that in this general formulation the specific orthogonality of $V$ directly leads to constraints in optimization, such as a system of nonlinear equations for the entries of $V$, which would need to be dealt with in the optimization routine. Of course, orthogonality of $V$ can be also enforced by the construction of the matrix, such as a rotation matrix and utilize the angles as degrees of freedom as demonstrated in the results section 3.3.
@@
$\text{arg max}_{\,(V,\vec{w})}\;\frac{1}{N\,R}\sum\limits_{i=1}^{N}\sum\limits_{r=1}^{R}\prod\limits_{l=1}^{3}g_{l}\left(p_{2}\left(V^{T}\cdot(\vec{z}_{i,r}+\vec{w});\vec{\gamma}_{l}\right)\right)\;,$
subject to $V\in\mathbb{R}^{3\times 3}$ with orthonormal columns and with $\vec{z}_{i,r}$ sampled from the densities $q_{i}(\vec{z})$.
A short comment regarding the calibration of this ranging/lateration system: The three lateration devices have specified coordinates $\vec{\gamma}_{1},\vec{\gamma}_{2},\vec{\gamma}_{3}$. Since the absolute position of the system in a global coordinate system is not important for practical measurements with that system, one can set $\vec{\gamma}_{1}=(0,0,0)^{T}$, $\vec{\gamma}_{2}=(c_{2,1},0,0)^{T}$ and $\vec{\gamma}_{3}=(c_{3,1},c_{3,2},0)^{T}$ and, consequently, obtain pose estimations relative to this defined geometry.